Find the equation of the tangent plane of the following surface patches
at the indicated points:
(b) σ(r, θ) = (r cos θ, r sin θ, 2θ), P =

3, 1,
π
3
1
Expert's answer
2020-10-26T19:12:45-0400

Solution:

Given,

"\sigma:\R^2\rightarrow \R^3\\
(r,\theta)\mapsto(r\cos \theta,r\sin \theta, 2\theta)"

We have to find the tangent plane of "\sigma" at "P(\sqrt{3}, 1,\pi/3)=\sigma(2,\pi/6)"

Now,

We have that

"\frac{\partial \sigma}{\partial r}=\sigma_r=(\cos\theta,\sin\theta, 0)""\frac{\partial \sigma}{\partial \theta}=\sigma_\theta=(-r\sin\theta, r\cos\theta,2)"


Thus,

"\sigma_r(2,\pi/6)=(\sqrt{3}/2,1/2,0)"


and

"\sigma_\theta(2,\pi/6)=(-1,\sqrt{3},2)"


Now we find the normal to the plane


"\vec n=\sigma_r(2,\pi/6)\times \sigma_\theta(2,\pi/6)=\begin{vmatrix}i &j&k
\\ \sqrt{3}/2&1/2&0\\
-1&\sqrt{3}&2
\end{vmatrix}=(1,-\sqrt{3},2)"



Hence,the equation of the tangent plane at point "P" is:


"1(x-\sqrt{3})-\sqrt{3}(y-1)+2(z-\pi/3)=0\\
\implies x-\sqrt{3}y+2z-2\pi/3=0"